The country of Albanystan (country A) forms a small open economy on its own. It has savings S_A=30+300r and investment I_A=80-200r. If r=0.2, is country A a borrower or a lender? Below what interest rate would country A be a borrower? The country of Bostonland (country B) decides to join and together they form a large open economy with S_B=40+100r and I_B=5-500r What is the equilibrium interest rate? What is CA_A,CA_B? Which country borrows and which country lend?
Answer #1

A) r=0.2 Savings=30+300*0.2=90 And investment = 80-200*0.2=40 Hence savings > investment Hence the country has enough savingsb. Now, plugging the value of the equilibrium interest rate into the national savings function of country B, we can get: SB=4

Kindly Upvote please